# A Comprehensive Survey of Advanced Persistent Threat Attribution: Taxonomy, Methods, Challenges and Open Research Problems

### v1

@@ Abstract
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
When firewalls and antivirus softwares become mere hurdles rather than barriers, we enter the realm of Advanced Persistent Threats (APTs). APT attacks are attack campaigns orchestrated by highly organized and often state-sponsored threat groups that operate covertly and methodically over prolonged periods [12]. APTs set themselves apart from conventional cyber-attacks by their stealthiness, persistence, and precision in targeting [1]. Unlike typical cyber-attacks, which aim for immediate and indiscriminate disruption or data exfiltration, APTs are orchestrated by resourceful adversaries who carefully conduct multi-phase operations. These adversaries excel at evading detection over prolonged periods, continuously adapting to security measures. APTs are recognized for their underlying intent and the sophisticated attack methodology [1]. APTs employ meticulously crafted tactics, techniques, and procedures (TTPs) tailored to breach specific targets or organization, in contrast to traditional attacks that often rely on automated tools with widespread and indiscriminate targets. This high degree of customization and persistence makes APTs exceptionally difficult to detect and counteract.
@@
Paper Year Attribution Case Study Artifact Taxonomy Method Classification Dataset Classification Challenges Open Research Problems
[88] $2022$
[98] $2022$
[43] $2023$
[42] * $2024$
Our Survey $2024$
    This survey is specific to malware-based attribution methods
Attributing attacks involves a multifaceted approach, combining digital forensics, intelligence analysis, and geopolitical insights. These approaches peel back the layers of anonymity or evasiveness that attackers hide behind [36, 61]. This intricate process employs a blend of technical and non-technical artifacts to identify the unique digital fingerprints left by attackers. Threat analyst teams scrutinize these indicators, comparing them against vast databases of known threats. It further aids in finding similarities or establishing new profiles for unidentified attackers. Analysts perform this analysis on traditional frameworks such as Diamond Model [3], CAM [24], Q model [17], triangle model [6], and MICTIC framework [7], which facilitate multi-dimension approach to connect the dots between collected artifacts and suspected threat actor. While traditional manual methods allow for deep and contextual analysis of attacks, their effectiveness and scalability are increasingly challenged by the rapidly evolving cyber threat landscape [88, 17]. The advent of advanced automated technologies such as artificial intelligence (AI) and machine learning (ML) presents an opportunity to enhance APT attribution’s performance, efficiency, reliability, and scalability. Further, the literature of automated threat attribution leverages series of attack artifacts with state-of-the-art automated technological capabilities to uncover the threat actors. We review such available automated threat attribution methods and present a comprehensive survey of automated APT attribution.
@@ I-B Review of prior surveys and gap analysis
The current surveys and research papers have explored various aspects of APTs and attribution techniques. Previous surveys have predominantly focused on traditional attribution methods [88, 43], the detection and mitigation of APTs [1, 43, 56], profiling techniques [98], threat intelligence of APTs [9, 62, 2], and threat modelling of APTs [20]. While these surveys provide valuable insights about APTs, there is a notable gap in the literature concerning the comprehensive review of automated APT attribution methods. The literature on the APT attribution methods is dispersed and scattered. In our exploration, we find very few surveys present in the literature [88, 98, 43, 42], which discuss threat attribution, but that too lacks detailed discussion, comparison of available attribution methods and current attribution challenges. There is also a lack of systematic literature that can summarize the complete flow and end-to-end automated attribution process. Such summarization facilitates a broader understanding of the community. Therefore, this paper systematizes the automated APT attribution methods available in the literature. The comprehensive survey aids in understanding the available effective automated attribution methodology. Table I compares this survey with the related existing study.
@@ I-C Scope and contribution
    We identify and propose potential open research problems in automated attribution, suggesting areas where further investigation and innovation can significantly advance threat attribution.
@@ I-D Structure of the Paper
The structure of this paper is shown in Fig 1. Section II presents the current state of APT attribution, and Section III details an attribution case study of past attack campaigns. Then, the proposed attribution artifact taxonomy is presented in Section IV, and a survey of the current attribution datasets and methods is discussed in Section V. Further, we present the current attribution challenges in Section VI and discuss open research problems and future research directions in Section VII. Section VIII concludes the presented survey. The list of acronyms used in this paper is listed in Table II.
@@ III-A SolarWinds cyberattack
    Operational TTPs: The attackers’ operational tactics also played a vital role in the attribution process. In the Solarwinds attack, the attackers demonstrated these advanced tactics by maintaining a low profile within the networks for several months, exfiltrating data without triggering alarms [23].
@@
The SolarWinds campaign showcases the complexity and sophistication of advanced cyber espionage campaigns. By examining artifacts such as malware analysis, C&C infrastructure, code overlaps, operational tactics, and victimology, cybersecurity experts attributed the campaign to APT29 (also known as Cozy Bear or Nobelium) [44, 129]. The reverse engineering revealed sophisticated coding, obfuscation techniques, and code similarities between the SolarWinds malware and previous APT29 campaigns [60, 125], providing key evidence for attribution. The use of DGA, randomization techniques, and the long dwell time were consistent with APT29’s past behavior [44]. The targeted intelligence-gathering aligned with APT29’s strategic objectives, further supporting the attribution of the SolarWinds espionage campaign to the Russia-backed group APT29[113, 136]. The convergence of these multiple evidence provided a robust foundation for attribution. This analysis highlights the importance of comprehensive and multi-layered analysis in identifying and understanding sophisticated threat actors.
@@ IV Systemization of Attribution Artifacts
Attribution is a multifaceted approach and is decided based on multiple clues. We call these clues as attribution artifacts. We define attribution artifacts as any critical pieces of evidence or indicators or information which aid in attributing the cyber threats, i.e., linking cyber threats to specific threat actors or origins or sources. during Russia’s invasion of Ukraine, the AcidRain wiper malware was employed in the Viasat attack [89, 101], and further analysis uncovered code overlaps between AcidRain wiper and the SolarWinds malware [44].
@@ IV-A Attribution artifacts taxonomy
Artifacts plays a crucial role in the attribution process. It is essential to understand the various types that can help security analysts link or profile the threat actor. Such understanding enhances defenders’ ability to trace attack origins and identify the techniques used. In this section, we present the taxonomy of artifacts as shown in Fig. 4 Our taxonomy of attribution artifacts categorizes the diverse information that can be utilized in the cyber threat attribution process. Based on source of the attribution artifacts, we categorize them in two types: 1) Attack artifacts 2) Non-attack artifacts. The details about these artifact categories and further classification are discussed in this section.
@@ IV-A1 Attack artifacts
It consists the artifacts which analysts captured from attacked infrastructure such as malware, IoCs, and toolchains etc. Based on nature of the artifacts, we categorize attack artifacts into two category: 1) Evidentiary artifacts and 2) Behavioral artifacts. The details about these artifact categories and further classification are discussed below.
@@
    Behavioral Artifacts: The methods attackers use, including their operational patterns and decision-making processes, are known as behavioral artifacts. Understanding these behaviors can reveal the attackers’ level of sophistication, potential motives, and operational goals. It is crucial to identify the likely perpetrators behind cyber attacks. We further divide behavioral artifact into five types as explained below:
@@
      Tactics Techniques and Procedures (TTPs): It describes how threat actors operate and offers valuable insights into their behavior and capabilities. Specific TTPs can act as signatures for certain groups, revealing their unique modus operandi such as using phishing to perform initial access or employing PowerShell for execution. The TTPs can be extracted from historical threat reports [81, 100, 90] or threat intelligence feeds to create the dataset. The TTP extraction methods help attackers to achieve their objectives and provide a detailed picture of their operational approach. The MITRE ATT&CK framework includes $12$ different type of tactics as shown in Fig 4.
@@
      Malware & Code analysis: The artifacts related to malware analysis and the source code of malware are crucial for understanding its functionality, propagation methods, and origins. Unique code snippets, code reuse, or targeted vulnerabilities can link malware to known threat actors [18]. For example, unique malware functions and reusing specific code blocks seen in other known malware can provide valuable insights. The capabilities of the malwares can also be extracted and correlate them to identify similar ones [124, 137]. The process of dissecting malware helps identify its behavior and origin. It involves performing static, dynamic, and hybrid analysis of malware samples.
@@ IV-A2 Non-attack Artifacts
It consists the artifacts which analysts captured from outside of the attacked infrastructure such as threat intelligence feeds, public claims, and geopolitical tension etc. Based on nature of the artifacts, we categorize non-attack artifacts into two category: 1) Open Source INTelligence (OSINT) artifacts and 2) External artifacts. The details about these artifact categories and further classification are discussed below.
@@
    Open Source INTelligence (OSINT) Artifacts: It includes the artifacts derived from publicly available sources, such as social media, threat reports, technical databases of historical attacks, and hacker forums. These artifacts can provide contextual and corroborative information to support attribution by linking public disclosures or activities to specific cyber incidents or threat actors. OSINT artifacts offer valuable insights into the threat actor’s activities, intentions, and associations. We further divide this artifact into three types, explained below:
@@
      Public Claims and Hacker Forms: The online platforms, including hacker forums, dark websites, and social media channels, are crucial sources of information where cybercriminals discuss tactics, share tools, and claim responsibility for cyber incidents. Although these claims need verification, they can provide direct links to the perpetrators. For instance, threat group blogs on dark websites, telegram channels, and social media posts by groups claiming an attack are crucial sources of information [132, 95]. Information or tools shared on public platforms can often be directly linked to specific cyber incidents or groups, such as a thread in a dark web forum where an actor sells access credentials to compromised systems.
@@
      Threat Intelligence Databases: The databases aggregate and analyze IoCs, attack patterns, and threat actor profiles. By correlating data from historical threat databases with indicators captured during an incident, analysts can establish links and identify related known threat actors. These repositories of threat data are invaluable for identifying and understanding cyber threats. For example, STIX/TAXII feeds and proprietary threat intelligence platforms.
@@
      Geopolitical: Information related to international relations, conflicts, and political events that may motivate or be exploited by cyber threat actors is crucial. We can identify possible state sponsorship or politically motivated actors by correlating cyber attacks with geopolitical events [7]. A rise in cyber espionage targeting government agencies during periods of heightened diplomatic tensions between two countries can indicate geopolitical motives. We witnessed such campaigns during recent Russia-Ukraine conflicts [37, 57].
@@ IV-B Comparison of automated attribution artifacts
Relevance measures how directly an artifact contributes to the identification of a specific threat actor. High relevance artifacts, such as IOCs, TTPs, C&C infrastructure, and malware & code analysis, provide strong and often unique identifiers that can clearly tie activities to specific threat actors, making them crucial in the attribution process. Medium relevance artifacts, like toolchains, public claims, threat intelligence database, geopolitical, and time zone analysis, offer valuable context but may require additional corroboration to be definitive, reflecting their role in supporting rather than leading attribution efforts. Low relevance artifacts, such as cryptocurrency transactions, language and writing styles, victimology, and post incident communication, provide useful supplementary information but often lack the direct connection needed for strong attribution, making them less central to the process. This analysis aligns with the idea that artifacts vary in their directness and specificity when attributing cyber threats.
@@
Credibility refers to the reliability of the source from which an artifact is obtained. It measures how likely it is that the information provided by the source is accurate and unbiased. High credible artifacts, such as those providing TTPs, C&C infrastructure data, TTPs, toolchains, and malware analysis, are typically derived from detailed investigations by reputable cybersecurity professionals and organizations, ensuring the reliability and accuracy of the information. Medium credible artifacts, such as geopolitical analysis, and threat intelligence dataases, offer useful insights but may come from a mix of reliable and less verified sources, requiring cross-verification to ensure accuracy. Low credible artifacts, such as public claims and hacker forums, are often unverified and can be intentionally misleading, making them less trustworthy for confident attribution. This analysis emphasizes the need to prioritize artifacts from high-veracity sources to ensure accurate and credible cyber threat attribution.
Timeliness measures how quickly an artifact can be obtained and analyzed after an attack. High timeliness artifacts, such as IOCs, toolchains, and time zone analysis, are often available almost immediately during or after an attack, allowing for rapid response and mitigation. Medium timeliness artifacts, like C&C infrastructure, TTPs, victimology, public claims, and malware & code analysis typically require some time for monitoring, analysis, and pattern recognition, which may delay immediate action but still offer relatively prompt insights. Low timeliness artifacts, such as cryptocurrency transactions, language & writing styles, threat intelligence databases, post incident communications and geopolitical considerations, involve more extensive investigation and slower information gathering, making them less suited for real-time threat response. This analysis emphasizes the importance of prioritizing timely artifacts in situations where rapid attribution and response are critical.
@@
This comparative analysis enhances the reliability of attribution, enabling more precise identification of adversaries. A glimpse of the reasoning behind this analysis is provided in Table XI in the Appendix. Moreover, understanding the strengths and limitations identified during comparison of each artifact type ensures a comprehensive approach, reduces the likelihood of false attributions, and improves overall threat intelligence.

### v3

@@ I Introduction
When firewalls and antivirus softwares become mere hurdles rather than barriers, we enter the realm of Advanced Persistent Threats (APTs). APT attacks are attack campaigns orchestrated by highly organized and often state-sponsored threat groups that operate covertly and methodically over prolonged periods [12]. APTs set themselves apart from conventional cyber-attacks by their stealthiness, persistence, and precision in targeting [50, 1]. Unlike typical cyber-attacks, which aim for immediate and indiscriminate disruption or data exfiltration, APTs are orchestrated by resourceful adversaries who carefully conduct multi-phase operations. These adversaries excel at evading detection over prolonged periods, continuously adapting to security measures. APTs are recognized for their underlying intent and the sophisticated attack methodology [1]. APTs employ meticulously crafted tactics, techniques, and procedures (TTPs) tailored to breach specific targets or organization, in contrast to traditional attacks that often rely on automated tools with widespread and indiscriminate targets. This high degree of customization and persistence makes APTs exceptionally difficult to detect and counteract.
@@
Paper Year Attribution Evolution Attribution Case Study Artifact Taxonomy Method Classification Dataset Classification Challenges Open Research Problems
[88] $2022$  ✗
[98] $2022$  ✗
[43] $2023$  ✗
[42] * $2024$  ✓
Our Survey $2024$  ✓
    This survey is specific to malware-based attribution methods only
Attributing attacks involves a multifaceted approach, combining digital forensics, intelligence analysis, and geopolitical insights. These approaches peel back the layers of anonymity or evasiveness that attackers hide behind [36, 61]. This intricate process employs a blend of technical and non-technical artifacts to identify the unique digital fingerprints left by attackers [15]. Threat analyst teams scrutinize these indicators, comparing them against vast databases of known threats. It further aids in finding similarities or establishing new profiles for unidentified attackers. Analysts perform this analysis on traditional frameworks such as Diamond Model [3], CAM [24], Q model [17], triangle model [6], and MICTIC framework [7], which facilitate multi-dimension approach to connect the dots between collected artifacts and suspected threat actor. While traditional manual methods allow for deep and contextual analysis of attacks, their effectiveness and scalability are increasingly challenged by the rapidly evolving cyber threat landscape [88, 17]. The advent of advanced automated technologies such as artificial intelligence (AI) and machine learning (ML) presents an opportunity to enhance APT attribution’s performance, efficiency, reliability, and scalability [121]. Further, the literature of automated threat attribution leverages series of attack artifacts with state-of-the-art automated technological capabilities to uncover the threat actors. We review such available automated threat attribution methods and present a comprehensive survey of automated APT attribution.
@@ I-B Review of prior surveys and gap analysis
The current surveys and research papers have explored various aspects of APTs and attribution techniques. Previous surveys have predominantly focused on traditional attribution methods [88, 43], the detection and mitigation of APTs [1, 43, 56, 41], profiling techniques [98], threat intelligence of APTs [9, 62, 2], and threat modelling of APTs [20, 55]. While these surveys provide valuable insights about APTs, there is a notable gap in the literature concerning the comprehensive review of automated APT attribution methods. The literature on the APT attribution methods is dispersed and scattered. In our exploration, we find very few surveys present in the literature [88, 98, 43, 42], which discuss threat attribution, but that too lacks detailed discussion, comparison of available attribution methods, discussion of current attribution challenges and open research problems. There is also a lack of systematic literature that can summarize the complete flow and end-to-end automated attribution process. Such summarization facilitates a broader understanding of the community. Therefore, this paper systematizes the automated APT attribution methods available in the literature. The comprehensive survey aids in understanding the available effective automated attribution methodology. Table I compares this survey with the related existing study.
@@ I-C Scope and contribution
    We identify and propose potential open research problems in automated attribution, suggesting novel problems and areas where further investigation and innovation can significantly advance automated threat attribution.
@@ I-D Structure of the Paper
The structure of this paper is shown in Fig 1. Section II presents the evolution and current state of APT attribution, and Section III details an attribution case study of past attack campaigns. Then, the proposed attribution artifact taxonomy is presented in Section IV, and a survey of the current attribution datasets and methods is discussed in Section V. Further, we present the current attribution challenges in Section VI and discuss open research problems and future research directions in Section VII. Section VIII concludes the presented survey. The list of acronyms used in this paper is listed in Table II.
@@ III-A SolarWinds cyberattack
    Operational TTPs: The attackers’ operational tactics also played a vital role in the attribution process. In the Solarwinds attack, the attackers demonstrated these advanced tactics by maintaining a low profile within the networks for several months, using legitimate credentials, and exfiltrating data without triggering alarms [97, 23].
@@
The SolarWinds campaign showcases the complexity and sophistication of advanced cyber espionage campaigns. By examining artifacts such as malware analysis, C&C infrastructure, code overlaps, operational tactics, and victimology, cybersecurity experts attributed the campaign to APT29 (also known as Cozy Bear or Nobelium) [44, 129]. The reverse engineering revealed sophisticated coding, obfuscation techniques, and code similarities between the SolarWinds malware and previous APT29 campaigns [60, 125], providing key evidence for attribution. The use of DGA, randomization techniques, and the long dwell time were consistent with APT29’s past behavior [44]. The targeted intelligence-gathering aligned with APT29’s strategic objectives, further supporting the attribution of the SolarWinds espionage campaign to the Russia-backed group APT29[113, 136]. The convergence of these multiple evidence provided a robust foundation for attribution. This analysis highlights the importance of comprehensive and multi-layered artifact analysis in identifying and understanding sophisticated threat actors.
@@ IV Systemization of Attribution Artifacts
Attribution is a multifaceted approach and done based on multiple clues. We call these clues as attribution artifacts. We define attribution artifacts as any critical pieces of evidence or indicators or information which aid in attributing the cyber threats, i.e., linking cyber threats to specific threat actors or origins or sources. During Russia’s invasion of Ukraine, the AcidRain wiper malware was employed in the Viasat attack [89, 101], and further analysis uncovered code overlaps between AcidRain wiper and the SolarWinds malware [44].
@@ IV-A Attribution artifacts taxonomy
Artifacts play a crucial role in the attribution process. It is essential to understand the various artifact types that can help security analysts to link or profile the threat actor. Such understanding enhances defenders’ ability to trace attack origins and identify the techniques used. In this section, we present the taxonomy of artifacts as shown in Fig. 4. Our taxonomy of attribution artifacts categorizes the diverse information that can be utilized in the cyber threat attribution process. Based on source of the attribution artifacts, we categorize them in two types: 1) Attack artifacts 2) Non-attack artifacts. The details about these artifact categories and further classification are discussed in this section.
@@ IV-A1 Attack artifacts
It consists the artifacts which analysts captured from compromised infrastructure such as malware, IoCs, and toolchains etc. Based on nature of the artifacts, we categorize attack artifacts into two category: 1) Evidentiary artifacts and 2) Behavioral artifacts. The details about these artifact categories and further classification are discussed below.
@@
    Behavioral artifacts: The methods attackers use, including their operational patterns and decision-making processes, are known as behavioral artifacts. Understanding these behaviors can reveal the attackers’ level of sophistication, potential motives, and operational goals. It is crucial to identify the likely perpetrators behind cyber attacks. We further divide behavioral artifact into four types as explained below:
@@
      Tactics Techniques and Procedures (TTPs): It describes how threat actors operate and offers valuable insights into their behavior and capabilities [46]. Specific TTPs can act as signatures for certain groups, revealing their unique modus operandi such as using phishing to perform initial access or employing PowerShell for execution. The TTPs can be extracted from historical threat reports [81, 100, 90] or threat intelligence feeds to create the dataset. The TTP extraction methods help attackers to achieve their objectives and provide a detailed picture of their operational approach. The MITRE ATT&CK framework includes $12$ different type of tactics as shown in Fig 4.
@@
      Malware & code analysis: The artifacts related to malware analysis and the source code of malware are crucial for understanding its functionality, propagation methods, and origins. Unique code snippets, code reuse, or targeted vulnerabilities can link malware to known threat actors [18]. For example, unique malware functions and reusing specific code blocks seen in other known malware can provide valuable insights. The capabilities of the malwares can also be extracted and correlate them to identify similar ones [11, 124, 137]. The process of dissecting malware helps identify its behavior and origin. It involves performing static, dynamic, and hybrid analysis of malware samples.
@@ IV-A2 Non-attack artifacts
It consists the artifacts which analysts captured from outside of the compromised infrastructure such as threat intelligence feeds, public claims, and geopolitical tension etc. Based on nature of the artifacts, we categorize non-attack artifacts into two category: 1) Open Source INTelligence (OSINT) artifacts and 2) External artifacts. The details about these artifact categories and further classification are discussed below.
@@
    Open Source INTelligence (OSINT) artifacts: It includes the artifacts derived from publicly available sources, such as social media, threat reports, technical databases of historical attacks, and hacker forums. These artifacts can provide contextual and corroborative information to support attribution by linking public disclosures or activities to specific cyber incidents or threat actors [5]. OSINT artifacts offer valuable insights into the threat actor’s activities, intentions, and associations. We further divide this artifact into two types, explained below:
@@
      Public claims and hacker forums: The online platforms, including hacker forums, dark websites, and social media channels, are crucial sources of information where cybercriminals discuss tactics, share tools, and claim responsibility for cyber incidents. Although these claims need verification, they can provide links to the perpetrators. For instance, threat group blogs on dark websites, telegram channels, and social media posts by groups claiming an attack are crucial sources of information [132, 95]. Information or tools shared on public platforms can often be directly linked to specific cyber incidents or groups, such as a thread in a dark web forum where an actor sells access credentials of compromised systems.
@@
      Threat intelligence databases: The databases aggregate and analyze IoCs, attack patterns, and threat actor profiles. By correlating data from historical threat databases with indicators captured during an incident, analysts can establish links and identify related known threat actors [74, 102]. These repositories of threat data are invaluable for identifying and understanding cyber threats. For example, STIX/TAXII feeds and proprietary threat intelligence platforms.
@@
      Geopolitical: Information related to international relations, conflicts, and political events that may motivate or be exploited by cyber threat actors is crucial. We can identify possible state sponsorship or politically motivated actors by correlating cyber attacks with geopolitical events [7, 130]. A rise in cyber espionage targeting government agencies during periods of heightened diplomatic tensions between two countries can indicate geopolitical motives. We witnessed such campaigns during recent Russia-Ukraine conflicts [37, 57].
@@ IV-B Comparison of automated attribution artifacts
Relevance measures how directly an artifact contributes to the identification of a specific threat actor. High relevance artifacts, such as IoCs, TTPs, C&C infrastructure, and malware & code analysis, provide strong and often unique identifiers that can clearly tie activities to specific threat actors, making them crucial in the attribution process. Medium relevance artifacts, like toolchains, public claims, threat intelligence database, geopolitical, and time zone analysis, offer valuable context but may require additional corroboration to be definitive, reflecting their role in supporting rather than leading attribution efforts. Low relevance artifacts, such as cryptocurrency transactions, language and writing styles, and post incident communication, provide useful supplementary information but often lack the direct connection needed for strong attribution, making them less central to the process. This analysis aligns with the idea that artifacts vary in their directness and specificity when attributing cyber threats.
@@
Credibility refers to the reliability of the source from which an artifact is obtained. It measures how likely it is that the information provided by the source is accurate and unbiased. High credible artifacts, such as those providing TTPs, C&C infrastructure data, TTPs, toolchains, and malware analysis, are typically derived from detailed investigations by reputable cybersecurity professionals and organizations, ensuring the reliability and accuracy of the information. Medium credible artifacts, such as geopolitical analysis, and threat intelligence databases, offer useful insights but may come from a mix of reliable and less verified sources, requiring cross-verification to ensure accuracy. Low credible artifacts, such as public claims and hacker forums, are often unverified and can be intentionally misleading, making them less trustworthy for confident attribution. This analysis emphasizes the need to prioritize artifacts from high-veracity sources to ensure accurate and credible cyber threat attribution.
Timeliness measures how quickly an artifact can be obtained and analyzed after an attack. High timeliness artifacts, such as IoCs, and time zone analysis, are often available almost immediately during or after an attack, allowing for rapid response and mitigation. Medium timeliness artifacts, like C&C infrastructure, TTPs, victimology, public claims, and malware & code analysis typically require some time for monitoring, analysis, and pattern recognition, which may delay immediate action but still offer relatively prompt insights. Low timeliness artifacts, such as cryptocurrency transactions, language & writing styles, threat intelligence databases, post incident communications and geopolitical considerations, involve more extensive investigation and slower information gathering, making them less suited for real-time threat response. This analysis emphasizes the importance of prioritizing timely artifacts in situations where rapid attribution and response are critical.
@@
This comparative analysis enhances the reliability of attribution, enabling more precise identification of adversaries. A glimpse of the reasoning behind this analysis is also provided in Table X in the Appendix. Moreover, understanding the strengths and limitations identified during comparison of each artifact type ensures a comprehensive approach, reduces the likelihood of false attributions, and improves overall threat intelligence.